OpenStudy (anonymous):

Auburn uses her credit card to buy some clothes for $827.78. She can pay off up to $315 per month. The card has an annual rate of 23.3% compounded monthly. How much total interest will she pay? $30.69 $13.99 $37.04 $16.07

OpenStudy (anonymous):

you would figure out how many months it would take her to pay it off and then multiply the annual rate by however many months.

OpenStudy (anonymous):

how do I do that? please help

OpenStudy (anonymous):

well, 315 times two is not enough so by default you know it would be 3 months...

OpenStudy (anonymous):

you would then multiply the annual interest by three since it is per month

OpenStudy (anonymous):

so is it 37.04 then?

OpenStudy (anonymous):

i believe so.

OpenStudy (anonymous):

are you sure?

OpenStudy (anonymous):

well for the first month it would be just 827.78 times the annual interest. have @johnweldon1993 take a look at it

OpenStudy (johnweldon1993):

what i just did had NOTHING even close...so im not sure ...ill research it though...sorry...maybe someone else can help?

OpenStudy (anonymous):

ok well can you help me this one Elijah purchased a home for $123,700. The home appreciates about 5.5% each year. What is the value of the home after 13 years? $59,290.24 $131,276.59 $212,145.50 $248,114.23

OpenStudy (johnweldon1993):

so this will be Vfinal = Vintial(1+r)^t Vfinal = 123700(1+.055)^13 = 123700(1.055)^13 Vfinal = 248,114.23

OpenStudy (anonymous):

thanks so much! would you be willing to help me with a couple more or no?

OpenStudy (johnweldon1993):

i'm actually about to be getting off for the night...but i'm sure the rest of the intelligent people on here would be more than happy to help you :)

OpenStudy (anonymous):

ok thanks anyways:)